OpenStudy (anonymous):

tan^-1 (-1) find the exact value of the expression

OpenStudy (jdoe0001):

\(\bf tan^{-1}(-1)\implies tan(\theta) = -1\implies tan^{-1}(tan(\theta)) = tan^{-1}(-1)\\ \quad \\ \theta = tan^{-1}(-1)\)

OpenStudy (jdoe0001):

check your Unit Circle